Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 184     Entry time: 06 May 2023, 02:19
Author: Jan-Etienne Pudell 
Group: MID 
Subject: Shift summary 

- We get stable x-ray delivery with 130µJ.

 

- We could move the detector succesfully in and out of the LFoV configuration!

- We added a hole in in the black Kapton foil and increased the holes in the Kapton shielding of the jet. These were not positioned well, due to a misalignment of the ALAS. After this intervention we did not observe any scattering rings on the detector.

- We aligned the liquid jet and get a nice diffraction rings from water. 

ELOG V3.1.4-7c3fd00